Using traditional methods, it takes 8.38.3 hours to receive a basic driving license. A new license training method using Computer Aided Instruction (CAI) has been proposed. A researcher used the technique with 1212 students and observed that they had a mean of 7.97.9 hours with a variance of 3.243.24. A level of significance of 0.10.1 will be used to determine if the technique performs differently than the traditional method. Assume the population distribution is approximately normal. Find the value of the test statistic. Round your answer to three decimal places.
Answer #1

Ho :µ =8.3     

Ha :µ ╪8.3(Two tail test)

  

Level of Significance , α = 0.10

variance=3.24

sample std dev ,  s = √3.24 = 1.80

Sample Size ,n = 12

Sample Mean,  x̅ =7.9000

  

degree of freedom=DF=n-1=11

  

Standard Error , SE =    s/√n =1.8000/ √ 12=0.5196

t-test statistic= (x̅ - µ )/SE = (7.900-8.3) / 0.520    =-0.7698

  

p-Value=0.4576    [Excel formula =t.dist(t-stat,df) ]

Decision:p-value>α, Do not reject null hypothesis               

Conclusion:     There is not enough evidence to conlcude that technique performs differently than the traditional method. at 0.10 level of significance